# Towards Fixing Panic Bugs for Real-world Rust Programs

### v1

@@ Abstract.
The Rust programming language has garnered significant attention due to its robust safety features and memory management capabilities. Despite its guaranteed memory safety, Rust programs still suffer from runtime errors that are unmanageable, i.e., panic errors. Notably, over half of the bugs in rustc, Rust’s own compiler, are attributable to crash stemming from panic errors. However, understanding root causes and resolving these panics often requires substantial effort due to the limited information provided, and the stack backtrace could be intricate, often omitting the actual fault locations. Although numerous automated program repair techniques exist, we observe that the prevailing fix patterns do not readily apply to Rust programs due to natural differences in language mechanisms.
To tackle the above challenges, this paper introduces a systematic study aimed at fixing Rust panic bugs. We commence by assembling a dataset, namely Panic4R, which includes 102 real panic bugs and their fixes from the top 500 most downloaded open-source crates. By analyzing Rust’s implementation, we identify Rust-specific patterns for fixing panic bugs, which can aid in understanding and providing guidance for generating patches. Finally, we design and implement the first automated fixing tool, PanicKiller, for Rust panic bugs, which effectively generates correct patches on the real-world large-scale dataset, and has already assisted in the resolution of 28 panic bugs in open-source projects. Each resolved issue has been validated by the developers and merged into the respective codebases.
Rust, automated program repair, fault localization
@@ 1. Introduction
As a statically typed programming language, Rust has gained popularity for its well-known memory safety guarantees and high performance. Recently, the White House Office of the National Cyber Director also emphasized the necessity of using programming languages that have fewer memory safety vulnerabilities (WhiteHou52:online, ), and nominating Rust as an example of a memory-safe programming language. The foundational principles of Rust, including ownership, borrowing, and lifetimes, enable developers to implement secure and efficient programs. Rust’s emphasis on zero-cost abstractions and fearless concurrency has significantly contributed to its popularity in systems programming (bugden2022rust, ; jung2021safe, ; klabnik2023rust, ; jung2020understanding, ). This focus has led to an increase in the development of widely recognized software projects written in Rust (Servothe7:online, ; RedoxYou24:online, ; TiKVTiKV64:online, ; StratisS70:online, ; Cloudfla15:online, ).
@@
Although Rust boasts security features and avoids common bugs such as null pointer dereference, uninitialized variables, and data races, it is still prone to panic errors, a Rust-specific crash type. The severity of the panic is evident from its typical consequences—program crashes or terminations that may lead to the improper handling of resources, such as unclosed file descriptors or network connections (Zheng2023A, ). Unlike Java’s structured exception handling framework that manages routine errors, Rust’s panic mechanism is designed for unrecoverable situations, significantly impacting program stability. Moreover, the Rust compiler, rustc, written in Rust, also exhibits vulnerabilities to panic bugs. Over half of the issues in Rust language’s official GitHub repository are categorized as Internal Compiler Errors (ICE) (Issues·r78:online, ), primarily caused by panic bugs in rustc. Therefore, understanding and resolving panic bugs in Rust is crucial for ensuring the reliability and stability of Rust programs.
However, fixing panic bugs can be a tedious and challenging task for Rust developers. On the one hand, the stack expansion information output by the compiler after a panic error is triggered could be complex and hard to understand. Regarding dozens of lines of stack backtrace, it is challenging to understand the root cause, localize the bug, and fix the program. Due to the unique memory management model and lifetime rules of Rust, the learning curve becomes particularly steep (Crichton2021The, ; crichton2024profiling, ; Qin2020Understanding, ). On the other hand, the infrastructure and toolchain supporting the Rust language are not as mature as other languages. Most of the existing code fix datasets or automated program repair (APR) tools are designed for other languages, such as Java and C/C++, but they could be difficult to migrate to Rust programs because of the differences in language mechanisms, which may result in the repaired program violating ownership rules or even failing to compile.
@@
Recently, a few program repairing tools have been proposed for Rust, yet they are insufficient to address the most severe panic-related bugs. Rust-lancet (yang2024lancet, ) was developed to tackle bugs related to violations of ownership rules through three specific strategies. However, these strategies are tailored exclusively to Rust’s ownership rules, which are verified prior to runtime, making them unsuitable for addressing panic bugs. Similarly, Ruxanne (robati2024patterns, ) and RustAssistant (deligiannis2023rustassistant, ) focus on common compilation issues, such as incorrect data types, but these patterns do not effectively mitigate panic bugs.
@@
To overcome the aforementioned challenges and fill the program understanding gap, in this paper, we design and implement the first infrastructure aiming at automatically fixing panic bugs for real-world Rust programs. We have constructed a dataset Panic4R, containing more than $100$ panic bug instances and corresponding fix patches, derived from open source projects in the ecosystem. Referring to the Rust implementation code, we further perform fix pattern mining with Rust-specific syntactic features to provide a reference for the community to better understand panic bugs and fixing strategies. Further, we implement an automated fixing tool for panic bugs, namely PanicKiller, which first applies dependency analysis for cross-file level error localization, then combines it with semantic information of the reported errors for fix pattern matching, and finally outputs sorted patches with scores and descriptions.
@@
  • Dataset. We have constructed the first public dataset for Rust panic bugs, named Panic4R, which comprises 102 real bugs and their corresponding fixes from PR records of the top 500 most downloaded Rust crates. Each bug-fix pair is meticulously organized and has undergone thorough manual verification, facilitating future research.
@@
The dataset, mined patterns, and the source code of PanicKiller can be found at https://sites.google.com/view/panickiller/home.
Figure 1. Overview of the proposed infrastructure.
2. Background and Motivation
@@
Rust, a statically typed programming language, is gaining popularity for its well-known memory safety guarantees and high performance, making it notably distinct from C/C++ and JVM family of languages. The memory and thread safety is ensured through a specific mechanism known as the ownership system, which naturally prevents common errors like null pointer errors. Consider the code snippet presented in Listing 2, where line 2 declares a variable named name of type Option<String>. This type encapsulates a String object and employs Option, a Rust-specific data structure designed to handle potentially null values. It uses Some to indicate the presence of a specific value and None to represent the absence of any value, as demonstrated in line 3. In the subsequent match expression, the scenarios of Some and None for name are addressed separately. If name contains a value, this value is assigned to the variable n, which is then utilized to construct a new string. During this process, the ownership of n is transferred from name, rendering name subsequently null. As a result, in line 5, the ownership of name has been transferred to greeting via the match expression. Consequently, name is now None. Attempting to unwrap it in line 8 will trigger a compile-time error due to the ownership transfer. Also, as shown in line 9, calling unwrap() without checking the null value will result in a runtime error, i.e., panic error, and the program terminates because it cannot handle the exception.
{listing}
[h] {minted}[linenos=false, fontsize=, highlightcolor=gray!10, highlightlines=5,8,9]rust 1 fn main() 2 let name: Option¡String¿ = Some(String::from(”Alice”)); 3 let no_name: Option¡String¿ = None; 4 let greeting = match name 5 Some(n)=¿format!(”Hi, ”, n), //ownership transferred 6 None =¿ ”Hi, guest”.to_string(), 7 ; 8 // name.unwrap(); // compilation error 9 no_name.unwrap(); // PANIC 10 An example of Rust program.
@@
Motivation. The distinctive memory checking mechanism and stringent ownership transfer rules in Rust present significant challenges for understanding Rust programs and debugging them. Besides, if a complex program encounters a panic bug at runtime, much effort is required to perform error localization and determine the root cause.
Although there have been many APR techniques practiced on Java/C/C++, due to the unique design mechanics of Rust, most of the existing repair patterns are not applicable to Rust programs. Specifically, Java uses null to represent a variable that points to no object in memory. While convenient for indicating uninitialized states, this design often leads to problems, with most of Java’s fix patterns addressing null pointer exceptions (liu2019tbar, ; koyuncu2019ifixr, ; kim2013par, ; liu2019avatar, ; durieux2017nullpointer, ). For example, the function in Listing 2 tries to return a null value while the expected return type is int. It would trigger a very common null pointer bug, and the existing APR tool generally adds an extra condition to check if the value is null, then return a random data that fits the type requirement.
{listing}
[h] {minted}[linenos=false, fontsize=, highlightcolor=gray!10, highlightlines=3,4,5]java public static int test() Integer value = null; // Not allowed in Rust + if (value == null) // Exisiting fix pattern + return 0; + return value; // NullPointerException A minimized Java program triggering null pointer exception and the corresponding fix patch.
@@
However, conventional null pointer remediation strategies are largely ineffective in Rust due to the language’s rigorous handling of null values. As demonstrated in Listing 2, Rust does not employ a direct equivalent of null. Instead, it utilizes the Option<T> data type to explicitly manage scenarios where values might be absent. Data that could be null must be encapsulated within an Option<T>, effectively preventing null pointer errors inherent to Rust’s design. Consequently, traditional fix patterns that address null pointer issues do not apply to Rust programs. Besides, we notice that the C/C++’s bug fixing dataset, CVE-Fixes (fu2022vulrepair, ; chen2023vrrepair, ), covers a wide variety of memory bugs such as improper restriction of operations within the bounds of a memory buffer (CWE-119), out-of-bounds write (CWE-787), and null pointer dereference (CWE-476), etc. However, these specific bug types are inherently untriggerable in safe Rust programs due to the language’s stringent safety guarantees. Thus, the methods commonly applied in C/C++ contexts are not transferable to Rust programs.
3. Infrastructure
@@
The workflow and architecture are illustrated in Figure 1. Initially, we gather real-world panic bugs and their corresponding patches from the ecosystem to construct a dataset, Panic4R, the specifics of which are detailed in Section 3.1. Subsequently, referring to the implementation of Rust’s compiler rustc, and its standard libraries, we identify panic-fix patterns, including abstract patches and natural language descriptions, as elaborated in Section 3.2. Finally, we develop an automated fixing tool, PanicKiller, utilizing the extracted patterns. PanicKiller conducts dependency analysis on source programs, organizes patch priorities, and generates hybrid fixing suggestions, further explained in Section 3.3.
3.1. Dataset Construction
@@
To build a dataset containing real-world panic bugs and their patches, we follow the construction process of Defects4J (just2014defect4j, ), which is the most classic dataset in the APR task, and select the top $500$ downloaded crates from the public repository of Rust crates (cratesio6:online, ) for real bugs and fix patches extraction, so as to construct a real dataset. Specifically, we manually employ the following workflow to collect code and thus guarantee the quality and reliability of the dataset.
@@
1-50 36 25,882 13.60 syn, rand, regex, aho-corasick, num-traits, clap, serde_json, strsim, time, idna, hashbrown, proc-macro2, smallvec
51-100 35 25,988 13.58 percent-encoding, chrono, uuid, textwrap, nom, tokio, hyper, futures, toml
101-150 6 26,516 23.83 httparse, object, rustc-demangle, rustls, form_urlencoded, gimli
151-200 3 19,173 16.67 reqwest, num-bignit, rayon
201-250 6 3,616 9.70 bumpalo, filetime, fixedbitset, phf, prost
251-300 3 25,572 10.33 pest, serde-yaml, libm
301-350 2 44,882 10.50 prettyplease, bytemuck
351-400 2 1,853 12.00 cargo_metadata, tinytemplate
401-450 5 5,444 25.00 tar, plotters, pretty_assertions, yansi
451-500 4 28,760 23.25 crossbeam, brotli-decompressor, indicatif, md5
Total 102 21,805 15.87 51

### v2

# PanicFI: An Infrastructure for Fixing Panic Bugs in Real-World Rust Programs

@@ Abstract.
The Rust programming language has garnered significant attention due to its robust safety features and memory management capabilities. Despite its guaranteed memory safety, Rust programs suffer from runtime errors that are unmanageable, i.e., panic errors. Notably, traditional memory issues such as null pointer dereferences, which are prevalent in other languages, are less likely to be triggered in Rust due to its strict ownership rules. However, the unique nature of Rust’s panic bugs, which arise from the language’s stringent safety and ownership paradigms, presents a distinct challenge. Over half of the bugs in rustc, Rust’s own compiler, are attributable to crash stemming from panic errors. However, addressing Rust panic bugs is challenging and requires significant effort, as existing fix patterns are not directly applicable due to the design and feature of Rust language. Therefore, developing foundational infrastructure, including datasets, fixing patterns, and automated repair tools, is both critical and urgent.
This paper introduces a comprehensive infrastructure, namely PanicFI, aimed at providing supports for understanding Rust panic bugs and developing automated techniques. In PanicFI, we construct a dataset, Panic4R, comprising 102 real panic bugs and their fixes from the top 500 most-downloaded open-source crates. Then, through an analysis of the Rust compiler implementation , we identify Rust-specific patterns for fixing panic bugs, providing insights and guidance for generating patches. Moreover, we develop PanicKiller, the first automated tool for fixing Rust panic bugs, which has already contributed to the resolution of 28 panic bugs in open-source projects. The practicality and efficiency of PanicKiller confirm the effectiveness of the patterns mined within PanicFI. Furthermore, Panic4R serves as a benchmark for evaluating APR tools focused on Rust panic bugs. We believe the construction and release of PanicFI could enable the expansion of automated repair research tailored specifically to Rust programs, addressing unique challenges and contributing significantly to advancements in this field.
Rust, program repair, fault localization
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
As a statically typed programming language, Rust has gained popularity for its well-known memory safety guarantees and high performance. Recently, the White House Office of the National Cyber Director also emphasized the necessity of using programming languages that have fewer memory safety vulnerabilities (Whi, pair), and nominating Rust as an example of a memory-safe programming language. The foundational principles of Rust, including ownership, borrowing, and lifetimes, enable developers to implement secure and efficient programs. Rust’s emphasis on zero-cost abstractions and fearless concurrency has significantly contributed to its popularity in systems programming (Bugden and Alahmar, 2022; Jung et al., 2021; Klabnik and Nichols, 2023; Jung, 2020). This design has led to an increase in the development of widely recognized software projects written in Rust (Ser, 2023; Red, 2023; TiK, 2023; Str, 2023; Cloudflare, 2023).
@@
Although Rust boasts security features and significantly reduce common bugs such as null pointer dereference, uninitialized variables, and data races, it could suffer from panic errors, which are caused by a Rust-specific error-handling mechanism. The severity of the panic is evident from its typical consequences—program crashes or terminations that may lead to the improper handling of resources, such as unclosed file descriptors or network connections (Zheng et al., 2023). Unlike Java’s structured exception handling framework that manages routine errors, Rust’s panic mechanism is designed for unrecoverable situations, significantly impacting program stability. Moreover, the Rust compiler, rustc, written in Rust, also exhibits vulnerabilities to panic bugs. Over half of the issues in Rust language’s official GitHub repository are categorized as Internal Compiler Errors (ICE) (Iss, [n. d.]), primarily caused by panic bugs in rustc. Therefore, understanding and resolving panic bugs in Rust is crucial for ensuring the reliability and stability of Rust programs.
However, the infrastructure and toolchain supporting the Rust language are not yet as mature as those for other languages. Most existing code fix datasets and automated program repair (APR) tools are designed for languages like Java and C/C++. These tools can be challenging to adapt to Rust due to significant differences in language mechanisms, which may lead to violations of ownership rules or even failing to compile. For instance, commonly used strategies for fixing null pointer errors in Java are entirely inapplicable in Rust, as Rust’s language design inherently disallows null values. Similarly, repair patterns for memory and pointer errors in C/C++ are not transferrable to handling panic-related bugs in Rust, as Rust’s safety guarantees prevent such memory errors from occurring. In addition, Rust’s unique memory management model and lifetime rules leads to a steep learning curve, highlighting the urgent need for infrastructure to support panic bug fixes, particularly in complex, large-scale and real-world Rust programs.
@@
Due to the lack of mature infrastructure, such as datasets and repair patterns, fixing panic bugs can be a tedious and challenging task for Rust developers. Recently, a few program repairing tools have been proposed for Rust, yet they are insufficient to address the most severe panic-related bugs. Rust-lancet (Yang et al., 2024) was developed to tackle bugs related to violations of ownership rules through three specific strategies. However, these strategies are tailored exclusively to Rust’s ownership rules, which are verified prior to runtime, making them unsuitable for addressing panic bugs. Similarly, Ruxanne (Robati Shirzad and Lam, 2024) and RustAssistant (Deligiannis et al., 2023) focus on common compilation issues, such as incorrect data types, but these patterns do not effectively mitigate panic bugs.
@@
To overcome the aforementioned challenges and fill the program understanding gap, in this paper, we design and implement the first infrastructure PanicFI aiming at automatically fixing panic bugs for real-world Rust programs. We have constructed a dataset Panic4R, containing more than $100$ panic bug instances and corresponding fix patches, derived from open source projects in the ecosystem. Referring to the Rust implementation code, we further perform fix pattern mining with Rust-specific syntactic features to provide a reference for the community to better understand panic bugs and fixing strategies. Further, we implement an automated fixing tool for panic bugs, namely PanicKiller, which first applies dependency analysis for cross-file level error localization, then combines it with semantic information of the reported errors for fix pattern matching, and finally outputs sorted patches with scores and descriptions.
@@
  • Dataset. We constructed the first public dataset for Rust panic bugs, named Panic4R, which comprises 102 real bugs and their corresponding fixes from PR records of the top 500 most downloaded Rust crates. Each bug-fix pair is meticulously organized and has undergone thorough manual verification, facilitating future research.
@@
2. Preliminary and Motivation
Rust language leverages ownership mechanism to manage memory safety and thus naturally prevent common errors like null pointer errors. This language mechanism enforces strict rules at compile-time that manage memory usage and ownership, effectively eliminating common memory issues such as dangling pointers and data races. Rust language present panic mechanism for unrecoverable errors that signal bugs or critical conditions where continuing execution is not safe. This mechanism is different from the error handling approaches of many other languages, such as Java, C, and Python, which allow programs to catch errors and potentially recover from them.
Figure 1. An example of Rust program.
@@ 2. Preliminary and Motivation
For example, consider the code snippet presented in Figure 1, where line 2 declares a variable named name of type Option<String>. This type encapsulates a String object and employs Option, a Rust-specific data structure designed to handle potentially null values. It uses Some to indicate the presence of a specific value and None to represent the absence of any value, as demonstrated in line 3. In the subsequent match expression, the scenarios of Some and None for name are addressed separately. If name contains a value, this value is assigned to the variable n, which is then utilized to construct a new string. During this process, the ownership of n is transferred from name, rendering name subsequently null. As a result, in line 5, the ownership of name has been transferred to greeting via the match expression. Consequently, name is now None. Attempting to unwrap it in line 8 will trigger a compile-time error due to the ownership transfer. Also, as shown in line 9, calling unwrap() without checking the null value will result in a runtime error, i.e., panic error, and the program terminates because it cannot handle the exception.
@@
Motivation. The distinctive memory checking mechanism and stringent ownership transfer rules in Rust present significant challenges for understanding Rust programs and debugging them. Besides, if a complex program encounters a panic bug at runtime, much effort is required to perform error localization and determine the root cause. Although there have been many APR techniques practiced on Java/C/C++, due to the unique design mechanics of Rust, most of the existing repair patterns are not applicable to Rust programs. Specifically, Java uses null to represent a variable that points to no object in memory. While convenient for indicating uninitialized states, this design often leads to problems, with most of Java’s fix patterns addressing null pointer exceptions (Liu et al., 2019b; Koyuncu et al., 2019; Kim et al., 2013; Liu et al., 2019c; Durieux et al., 2017).
Figure 2. A minimized Java program triggering null pointer exception and the corresponding fix patch.
@@
For example, the function in Listing 2 tries to return a null value while the expected return type is int. It would trigger a very common null pointer bug, and the existing APR tool generally adds an extra condition to check if the value is null, then return a random data that fits the type requirement. However, conventional null pointer remediation strategies are largely ineffective in Rust due to the language’s rigorous handling of null values. As demonstrated in Figure 1, Rust does not employ a direct equivalent of null. Instead, it utilizes the Option<T> data type to explicitly manage scenarios where values might be absent. Data that could be null must be encapsulated within an Option<T>, effectively preventing null pointer errors inherent to Rust’s design. Consequently, traditional fix patterns that address null pointer issues do not apply to Rust programs. Besides, we notice that the C/C++’s bug fixing dataset, CVE-Fixes (Fu et al., 2022; Chen et al., 2023), covers a wide variety of memory bugs such as improper restriction of operations within the bounds of a memory buffer (CWE-119), out-of-bounds write (CWE-787), and null pointer dereference (CWE-476), etc. However, these specific bug types are inherently untriggerable in safe Rust programs due to the language’s stringent safety guarantees. Thus, the methods commonly applied in C/C++ contexts are not transferable to Rust programs.
@@
Our work. To address the aforementioned challenges, we construct a systematic infrastructure infra with dataset, fixing patterns and testing tools, the key components of which are illustrated in Figure 3. Initially, we gather real-world panic bugs and their corresponding patches from the ecosystem to construct a dataset, Panic4R, the specifics of which are detailed in Section 3. Subsequently, referring to the implementation of Rust’s compiler rustc, and its standard libraries, we identify panic-fix patterns, including abstract patches and natural language descriptions, as elaborated in Section 4. Finally, we develop an automated fixing tool, PanicKiller, utilizing the extracted patterns. PanicKiller conducts dependency analysis on source programs, organizes patch priorities, and generates hybrid fixing suggestions, further explained in Section 5.
Figure 3. Key components of the proposed infrastructure PanicFI.
3. Dataset Construction
@@
To build a dataset containing real-world panic bugs and their patches, we follow the construction process of Defects4J (Just et al., 2014), which is the most classic dataset in the APR task, and select the top $500$ downloaded crates from the public repository of Rust crates (cra, 2024) for real bugs and fix patches extraction, so as to construct a real dataset. Specifically, we manually employ the following workflow to collect code and thus guarantee the quality and reliability of the dataset.
@@
1-50 36 25,882 13.60 2137.2 79.90% syn, rand, regex, aho-corasick, num-traits, clap, serde_json, strsim, time, idna, hashbrown, proc-macro2, smallvec
51-100 35 25,988 13.58 409.2 86.50% percent-encoding, chrono, uuid, textwrap, nom, tokio, hyper, futures, toml
101-150 6 26,516 23.83 1578.6 90.20% httparse, object, rustc-demangle, rustls, form_urlencoded, gimli
151-200 3 19,173 16.67 216 85.50% reqwest, num-bignit, rayon
201-250 6 3,616 9.70 27.5 80.80% bumpalo, filetime, fixedbitset, phf, prost
251-300 3 25,572 10.33 78 63.80% pest, serde-yaml, libm
301-350 2 44,882 10.50 17 79.10% prettyplease, bytemuck
351-400 2 1,853 12.00 34.5 86.40% cargo_metadata, tinytemplate
401-450 5 5,444 25.00 58.7 80.70% tar, plotters, pretty_assertions, yansi
451-500 4 28,760 23.25 6 36.50% crossbeam, brotli-decompressor, indicatif, md5
Total 102 21,805 15.87 456.27 76.94% 51
4. Fix Pattern Mining
To comprehensively identify the causes and fix patterns for panic bugs in Rust programs, we analyze and infer data from the Rust implementation code (rus, [n. d.]). We avoid mining fix patterns solely from Panic4R to ensure diversity and comprehensiveness in the repair patterns. The root causes of panic errors in Panic4R may not be exhaustive; for instance, overflow issues may stem from different operators, such as addition or subtraction, which could be impossible to cover fully through real-world examples. In contrast, official implementation code includes all such cases and provides concise examples, making it more suitable for our analysis and collection.
In the Rust compiler source code, panic-related error messages are encapsulated within macros such as panic_const!(). These macros and their accompanying messages provide the diverse types of panics and their fundamental causes. We compile a summary of these categories and their descriptions in Table  2. Following this categorization, we examine specific scenarios within each category. Rust compiler developers mark each panic occurrence with the annotation # Panics, which details the causes and circumstances of the panic. These annotations are often paired with # Safety or # Examples, providing either fix strategies or bug-triggering examples. This annotated information allows us to delineate code-level fix patterns for each panic type.
Table 2. Root causes of panic bugs derived from Rust compiler source code.
Root Causes Code Examples
Unwrap on None/Invalid value
let x: Option<i32> = None;
let y = x.unwrap(); // PANIC
Mixed borrowing
let x = Rc::new(RefCell::new(5));
let borrow = x.borrow(); // immutable borrow
let borrow_mut = x.borrow_mut(); // PANIC
Async functions wrong resume async fn my_async_function(cx: &mut Context<’_>) {     let mut task = some_async_task(); // starts async task     let result = ready!(task.poll(cx)); // waits until ready     ready!(task.poll(cx)); // PANIC }
Arithmetic overflow
let x: i32 = i32::MAX; // maximum i32 value
x + 1; // PANIC
Index out of bounds
let array = [1, 2, 3, 4, 5];
array[5]; // PANIC
Invalid UTF-8 boundary
let s = "™"; // a UTF8 string (3 bytes)
&s[1..]; // PANIC
Division/Modulus by zero
let a = 10; let b = 0;
a / b; // PANIC
Assertion failed
assert!(1 == 0); // PANIC
Unreachable code
unreachable!(); // PANIC
Others
panic!("Here comes a panic!"); // PANIC
Figure 4. The implementation of borrow() function in RefCell, including annotation about potential panics.
For example, Figure 4 showcases a Rust source code snippet annotated to indicate a panic bug related to the ownership mechanism. The annotation includes an example where the panic is triggered by a double borrowing of a RefCell parameter, which is a Rust smart pointer facilitating internal mutability. To prevent such panics, the annotation suggests using the try_borrow() method as an alternative approach. By systematically reviewing all documented panic instances and their proposed resolutions in the Rust source code, we collect a comprehensive list of fix strategies.
To further delineate fix patterns, we convert code modifications into Abstract Syntax Trees (ASTs). By omitting irrelevant tokens, such as variable names, we identify similar AST transformation structures that reveal potential fixes for each identified root cause. In total, we identified 19 fix patterns, which include 34 sub-patterns involving different replacement strategies on one binary operator. A representative categorization along with corresponding examples is presented in Table 3. The complete list of patterns and sub-patterns is available on our website.